OpenStudy (anonymous):

Help here please: Are my answers correct?. b. Determine the measure of ∠ B if it exists. Answer: Using sine law for this q. sinB/7.2 = sin(35)/9.3 sin B = 7.2 * sin35/9.3 = 0.472 Angle(B) = sin-1(0.472) = 28.19 Therefore, Angle B = 28.19

OpenStudy (anonymous):

I already answer it, I just need to check whether my answers are correct

OpenStudy (anonymous):

@callisto:) hello again! Are they correct? I used formulas from my book

OpenStudy (asnaseer):

I'm not sure you've used the sine rule correctly here

OpenStudy (callisto):

a, correct b,i don't think so sinB/b = sin(A)/a sinB/9.3 = sin(35)/7.2 ....

OpenStudy (asnaseer):

\[\frac{9.3}{\sin(B)}=\frac{7.2}{\sin(35)}\]

OpenStudy (kinggeorge):

Now that callisto mentions it, I agree. My bad. :(

OpenStudy (anonymous):

Oh!, i got the orders mixed up. One sec, Il retry

OpenStudy (callisto):

Not really....

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

ok. Il review it one sec

OpenStudy (asnaseer):

hmmm - I get a different answer. and remember there should be two values - one acute and one obtuse

OpenStudy (callisto):

sinB = 9.3sin(35)/7.2 ....

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

I didnt add the ) to close it in my calculator thats why it mixed it up

OpenStudy (asnaseer):

I blame the fact that you're hanging up-side-down for that ;-)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok. so is the answer then 47.79?

OpenStudy (asnaseer):

close enough :)

OpenStudy (anonymous):

its not?

OpenStudy (asnaseer):

and the obtuse angle?

OpenStudy (callisto):

I've got 47.81 for the first angle. Remember you've the second possible angle

OpenStudy (asnaseer):

I got same as Callisto

OpenStudy (anonymous):

maybe i should have added more decimals:(. whats the formula for the obstute?

OpenStudy (asnaseer):

sin(x) = sin(180-x)

OpenStudy (anonymous):

ok so: is it 32.99?

OpenStudy (callisto):

the second angle is in quadrant II ....

OpenStudy (asnaseer):

no....

OpenStudy (callisto):

it should be greater than 90 degrees... it's an obtuse angle~

OpenStudy (anonymous):

Ok one second:)

OpenStudy (asnaseer):

you calculated sin(B) = 0.741... so when you take the inverse of this, you get two possible angles: x and 180-x

OpenStudy (asnaseer):

because sin(x) = sin(180-x)

OpenStudy (anonymous):

what is the value of x?

OpenStudy (asnaseer):

so call one angle x and the other one y and we have: y = 180 - x with: sin(x) = sin(y) = 0.741...

OpenStudy (callisto):

x is the angle you've just found

OpenStudy (asnaseer):

x=B in this case

OpenStudy (anonymous):

so is it then: 0.129?

OpenStudy (anonymous):

so its 7.41?

OpenStudy (asnaseer):

ok - slow down you understand the steps up to: sin(B) = 0.741...

OpenStudy (anonymous):

oh boy, this part im confused

OpenStudy (asnaseer):

sinB = 9.3sin(35)/7.2 = 0.741... agreed?

OpenStudy (anonymous):

Yes Agreed:)

OpenStudy (asnaseer):

ok, when you invert this to find B, you do: B = arcsin(0.741...) agreed?

OpenStudy (anonymous):

Agreed

OpenStudy (asnaseer):

that has many possible answers. but we are only interested in the first two quadrants. so the solutions are: B = 47.81 and (180 - 47.81)

OpenStudy (asnaseer):

understand?

OpenStudy (anonymous):

yes:)

OpenStudy (asnaseer):

phew! - we finally got there... :D

OpenStudy (anonymous):

:) THank you so much. sorry, Sometimes im really slow XD. Thank you for being patient and helping me:)

OpenStudy (asnaseer):

don't worry - at least you persevere which means you are learning. we've all been there!

OpenStudy (callisto):

^agree :)